Let 𝑓(π‘₯) = 1 3 π‘₯ 3 + π‘₯ 2 βˆ’ 15π‘₯ βˆ’ 9 . Use detailed sign tables in answering the following questions. (a) Find the intervals in which 𝑓 is increasing or decreasing. (b) Find the intervals in which the graph of 𝑦 = 𝑓(π‘₯) is concave upwards or downwards.Β 


1
Expert's answer
2021-04-29T16:54:55-0400

As I understood, the function is "f(x) = 13x^3+x^2-15x-9" .


a) Then, intervals in which f(x) is increasing "\leftrightarrow" "f'(x) > 0" .

Intervals in which f(x) is decreasing "\leftrightarrow""f'(x) < 0" .

"f'(x) = 39x^2+2x-15 \longrightarrow" find nulls of "f'(x)"

"x_{1,2} = \frac{1}{78}(-2 \pm \sqrt{4+4*15*39}) = \\
\frac{1}{39}(-1 \pm \sqrt{586}), x_1 < x_2"

We have found null of quadratic trinomial, so since it's "branches" are directed up :

"x < x_1 \longrightarrow f'(x) >0 \\
x_1 < x < x_2 \longrightarrow f'(x) <0 \\
x_2 < x \longrightarrow f'(x) >0"

So, the answer :

f(x) increases on "(-\infty; -\frac{1}{39} - \frac{\sqrt{586}}{39}) \: \cup \: (-\frac{1}{39} + \frac{\sqrt{586}}{39}; +\infty)"

f(x) decreases on "(-\frac{1}{39}-\frac{\sqrt{586}}{39};-\frac{1}{39} + \frac{\sqrt{586}}{39})"

b) Plot of "y = f(x)" is concave upwards if "f''(x) < 0" .

Plot of "y = f(x)" is concave downwards if "f''(x) > 0"

"f''(x) = (f'(x))' = 78x+2, f''(x_0) = 0 \longrightarrow x_0 = -\frac{1}{39}"

So, ... upwards if "(x < -\frac{1}{39})" , and ... downwards if "(-\frac{1}{39} < x)"